If masses 16 kg,  8kg and 4kg are in contact and a force of 140N applied on 4kg mass then what is force on  16 kg mass? 

Grade:12

2 Answers

kkbisht
5 years ago
when the  force of 140N acts on the combined mass of the blocks(16+8+4=28kg) the common acceleration of the blocks is 140/28=5m/s2
therefore the force on the 16Kg mass is 16x5=80N
 
kkbisht
Anshul Purandare
5 years ago
If masses are m1,m2&m3 and force be f
Then,
force on m1=(m1*f)/(m1+m2+m3)
              Consider m1= 16 kg
Then,
Force on 16 kg
              =16*140/(16+8+4)
              =16*140/28
              =80 N